Supplemental Operations Report – December 11th

Supplemental Operations Report 10-DEC-2022

Name of person filing report: Shannon Rupert
Reason for Report: Routine/crew changeover

Non-nominal systems: Nothing to report
Action taken for non-nominal systems: n/a
Generator: Running from about 8 pm to 8 am.
ScienceDome Dual Split: on at 65 degrees overnight

Solar— Nominal.
Solar— SOC Last 24 hours:
Max 100%
Min 59%
Avg 75.4%
Propane Reading, station tank – 55 %
Propane Reading, director tank— 57 %
Propane Reading, intern tank— 55 %
Propane Reading, generator— 61% (It’s pretty clear that the propane that was delivered last week was only to the generator tank, as it empties twice as fast as the other three.)
Ethanol Free Gasoline – 0 gallons
Water (static tank) – 550 gallons
Water in GreenHab – gallons
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 250 gallons
Hab toilet tank emptied: check regular operations reports

Sojourner rover used: no
Hours: not noted
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Notes on rovers: Oppy is no longer overheating

ATV’s Used: (Honda, 350.1, 350.2, 300): none
Reason for use: n/a
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Nothing to report
HabCar used and why, where? Yes, to town
CrewCar used and why, where? Yes, to town and to Grand Junction. Currently in Grand Junction for a week.
Luna used and why, where? Yes, to town
General notes and comments: We have started troubleshooting the power system issues.
Summary of internet: Nominal.
Summary of suits and radios: See crew operations reports.
Campus wide inspection, if action taken, what and why? Nothing to report
Summary of general operations: Conduit work delayed due to rain this week.

Summary of Hab operations: The crew reported that the humidifier on the upper deck was not working. We confirmed it wasn’t and in the process discovered that the humidifier was not actually one that we had purchased. Not really sure what happened there!

We fixed the door latch on the second stateroom door. This crew reported that they would have like some way to identify which door was theirs. Maybe we need to revisit the idea of numbers for the door.

We are set up and ready for the teachers of Spaceward Bound Utah Cohort 4 who arrive tomorrow.

Summary of Outpost operations: Sergii’s trailer now has carpet on the large tipout.

Summary of GreenHab operations: Plants were fertilized. Timer was installed but did not work. That was fixed today. I also realized that we still have the 90% shade cloth on so that will be removed tomorrow.

Summary of ScienceDome operations: Nothing to report.
Summary of RAM operations: Nothing to report
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nothing to report
Questions, concerns, supplies needed and requests: n/a

Copyright © The Mars Society. All rights reserved. | Main Site